OpenStudy (anonymous):

Find the exact value of cos 11π/3

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

\[\cos(11×\frac{ \pi }{ 3 })\] is it like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cos(2pi+5pi/3)

OpenStudy (anonymous):

OR \[\cos(\frac{ 11\pi }{ 3 })\]

OpenStudy (anonymous):

@SerikMB let the user understand dont just use https://www.wolframalpha.com/input/?i=Find+the+exact+value+of+cos+11%CF%80%2F3

OpenStudy (anonymous):

for the answer.

OpenStudy (anonymous):

@some_someone, he said that he knows how to find, i just wondering why then just don't use a calculator. xD

OpenStudy (anonymous):

in the previous question